OpenStudy (anonymous):

is x^2+4y^2=1 a function?

jimthompson5910 (jim_thompson5910):

Solve for y first: \[\large x^2+4y^2=1\] \[\large 4y^2=1-x^2\] \[\large y^2=\frac{1-x^2}{4}\] \[\large y=\pm\sqrt{\frac{1-x^2}{4}}\] So this means that \[\large y=\sqrt{\frac{1-x^2}{4}} \ \ \text{or} \ \ y=-\sqrt{\frac{1-x^2}{4}}\] But y can't be two different values for any given x value. So this is NOT a function.

OpenStudy (anonymous):

No because you have a y^2 term. Any equation with y^2 is not a function.

jimthompson5910 (jim_thompson5910):

that's one good way to remember it

OpenStudy (amistre64):

y^2 can be a function; .... its just that this equation is a circle, and therefore not a funtion